# Decoupled Classifier-Free Guidance for Counterfactual Diffusion Models

### v2

@@ F.1 Dataset details
F.2 Extra quantitative results for MIMIC-CXR
Table A.7: MIMIC: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{aff}}$. Compared to global CFG (i.e., $\omega$), DCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we apply $\omega_{\mathrm{inv}}=1.2$ to prevent degradation of invariant attributes. While reversibility tends to degrade as $\omega_{\mathrm{aff}}$ increases, DCFG maintains better reversibility than global CFG at higher guidance strengths.
do(key) CFG Sex AUC/$\Delta$ Race AUC/$\Delta$ Finding AUC/$\Delta$ MAE LPIPS
do(sex) $\omega=1.0$ 92.4 / +0.0 75.6 / +0.0 88.8 / +0.0 0.146 0.202
@@ F.2 Extra quantitative results for MIMIC-CXR
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 96.4 / +4.0 74.6 / -1.0 89.1 / +0.3 0.158 0.217
@@
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 95.4 / +0.3 69.6 / +4.2 90.3 / -0.1 0.141 0.198
@@
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 95.1 / +0.5 78.1 / -0.2 85.6 / +4.8 0.142 0.202
@@
Table A.8: MIMIC: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega=2.5$.
do(key) CFG Sex AUC/$\Delta$ Race AUC/$\Delta$ Finding AUC/$\Delta$ MAE LPIPS
do(sex) $\omega=1.0$ 92.4 / +0.0 75.6 / +0.0 88.8 / +0.0 0.146 0.202
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 99.9 / +7.5 71.3 / -4.3 86.2 / -2.6 0.200 0.261
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 91.3 / -3.8 89.5 / +24.1 88.4 / -2.0 0.181 0.237
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 93.0 / -1.6 77.0 / -1.3 99.0 / +18.2 0.143 0.206
@@
F.3 Extra visual results for MIMIC
@@
Figure A.15: Additional qualitative results for do(finding) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while accurately applying the intended intervention. Compared to standard CFG, DCFG produces counterfactuals with more localized changes and stronger identity preservation.
@@
Figure A.16: Additional qualitative results for do(sex) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while accurately applying the intended intervention on sex. Compared to standard CFG, which tends to amplify unrelated features such as disease (i.e. finding), DCFG produces counterfactuals with more localized, semantically aligned changes and stronger identity preservation.
@@
Figure A.17: Additional qualitative results for do(race) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). While race interventions correspond to relatively subtle visual changes, standard CFG often amplifies unrelated features such as disease appearance (e.g., finding). In contrast, DCFG better preserves invariant attributes and identity, producing counterfactuals that are more localized, semantically aligned, and faithful to the intended intervention.

### v5

@@ H.1 Dataset details
H.2 Extra quantitative results for MIMIC-CXR
Table A.11: MIMIC: Effectiveness (ROC-AUC $\uparrow$) and reversibility (MAE, LPIPS $\downarrow$) when varying $\omega_{\mathrm{aff}}$. Compared to global CFG, FCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we fix $\omega_{\mathrm{inv}}{=}1.2$ to prevent degradation of invariant attributes.
[…]
do(sex) CFG (Ho and Salimans, 2022) $\omega=1.0$ 92.4 / +0.0 75.6 / +0.0 88.8 / +0.0 0.146 0.202
@@ H.2 Extra quantitative results for MIMIC-CXR
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 96.4 / +4.0 74.6 / -1.0 89.1 / +0.3 0.158 0.217
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 95.4 / +0.3 69.6 / +4.2 90.3 / -0.1 0.141 0.198
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 95.1 / +0.5 78.1 / -0.2 85.6 / +4.8 0.142 0.202
@@
Table A.12: MIMIC: Effectiveness (ROC-AUC $\uparrow$) and reversibility (MAE, LPIPS $\downarrow$) when varying $\omega_{\mathrm{inv}}$. Increasing $\omega_{\mathrm{inv}}$ increases effectiveness on invariant variables while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, amplification on invariant attributes becomes comparable to the global CFG setting with $\omega=2.5$.
[…]
do(sex) CFG (Ho and Salimans, 2022) $\omega=1.0$ 92.4 / +0.0 75.6 / +0.0 88.8 / +0.0 0.146 0.202
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 99.9 / +7.5 71.3 / -4.3 86.2 / -2.6 0.200 0.261
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 91.3 / -3.8 89.5 / +24.1 88.4 / -2.0 0.181 0.237
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 93.0 / -1.6 77.0 / -1.3 99.0 / +18.2 0.143 0.206
@@
H.3 Extra visual results for MIMIC
@@
Figure A.15: Additional qualitative results for do(finding) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). FCFG better preserves invariant attributes and identity while accurately applying the intended intervention. Compared to standard CFG, FCFG produces counterfactuals with more localized changes and stronger identity preservation.
@@
Figure A.16: Additional qualitative results for do(sex) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). FCFG better preserves invariant attributes and identity while accurately applying the intended intervention on sex. Compared to standard CFG, which tends to amplify unrelated features such as disease (i.e. finding), FCFG produces counterfactuals with more localized, semantically aligned changes and stronger identity preservation.
@@
Figure A.17: Additional qualitative results for do(race) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). While race interventions correspond to relatively subtle visual changes, standard CFG often amplifies unrelated features such as disease appearance (e.g., finding). In contrast, FCFG better preserves invariant attributes and identity, producing counterfactuals that are more localized, semantically aligned, and faithful to the intended intervention.